Circular No. 8719

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 L1 (GARRADD) Gordon J. Garradd reports his discovery of a diffuse comet on images

taken with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation below). Follow-up observations by R. H. McNaught on June 6.4 UT show a 15′′ tail in p.a. 150^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
June 4.37559  10^{h}02^{m}22.^{s}77  −34^{o}18′56′′.0  17.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-L35.

T = 2006 Oct. 17.242 TT  ω = 336.^{o}615
Ω = 101.716 }2000.0
q = 1.49940 AU  i = 143.222

COMETS C/2006 G3, C/2006 J1–J3, C/2006 K5 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8718) — all being Kreutz sungrazers except C/2006 K5 (no known group). C/2006 J1 was very small, slightly diffuse, and too faint for photometry. C/2006 J2 was small, stellar, and faint (mag ∼ 8.0). C/2006 J3 was slightly diffuse and reached mag 7.7 on May 4.496 at 6.4R_{⊙}; it showed a 162′′ tail on May 4.538 at 5.9R_{⊙}. C/2006 K5 was extremely faint, peaking at mag ∼ 8.5 at May 22.196, and measurer K. Battams could only find it in about 10 images. Corrigendum: C/2006 G3 was found by J. Sachs, not H. Su (cf. IAUC 8717).

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 J1 May 3.213 2^{h}44.^{m}6 +14^{o}13′ C2 HS 2006-L20
C/2006 J2  4.413 2 50.7 +14 43 C2 HS 2006-L20
C/2006 J3  4.479 2 51.5 +14 46 C2 TH 2006-L20
C/2006 K5  22.113 4 12.5 +22 43 C3 HS 2006-L20

CORRIGENDA On IAUC 8710, “RS Ophiuchi”, line 9, for Pashcen, read Paschen, On IAUC 8718, “Comet C/2006 HW_{51}”, line 1, for R. H. McNaught

read G. J. Garradd

2006 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green